OpenStudy (anonymous):

prove (sec^2x -1)/(sec^2x) = sin^2x

OpenStudy (anonymous):

sec(x)=1/cos(x) (1/cos^2(x)-1)/(1/cos^2(x))=(1/cos^2(x)-1)*cos^2(x)=1-cos^2(x)=sin^(x)

OpenStudy (aum):

\[ \frac{\sec^2(x) - 1}{\sec^2(x)} = \frac{\sec^2(x)}{\sec^2(x)} - \frac{1}{\sec^2(x)} = 1 - \cos^2(x) = \sin^2(x) \]

OpenStudy (anonymous):

thanks guys

OpenStudy (aum):

You are welcome.